OpenStudy (anonymous):

Really need help!! And check some answers. what is the simplified form of each expression. medals given check this one 1. n^6/n^2 =n^8 Check this one 2. c^8 d^-12/c^-4 d^-8= c^-4 d^4

OpenStudy (anonymous):

help with this one 3. \[\left( \frac{ 3x }{ 2 } \right)^{4}\]

OpenStudy (anonymous):

@ehuman Hey was wondering if you can help me out?

OpenStudy (dape):

You made a small error in the first one, you added the exponents instead of subtracting: \[\frac{n^6}{n^2}=\frac{nnnnnn}{nn}=nnnn=n^4\] You made similar errors in the second one, you should subtract what's in the denominator: \[\frac{c^8d^{-12}}{c^{-4}d^{-8}}=c^{8-(-4)}d^{-12-(-8)}=c^{12}d^{-4}\]

OpenStudy (dape):

For the third one you can use that \((xy)^n=x^ny^n\), or in words, when things are multiplied together and then are taken to a power you can instead first take each factor to that power and then multiply. So instead of doing (3 times x divided by 2) to the 4th power, we can take each of the stuff in parenthesis to the 4th power and then multiply/divide it: \[\left(\frac{3x}{2}\right)^4=\frac{3^4x^4}{2^4}=\frac{81x^4}{16}\]

OpenStudy (anonymous):

wow thank you so much! Could you help with two more possible?

OpenStudy (anonymous):

3. \[\left( \frac{ 2 }{ 5n ^{9} } \right)^{2}\]

OpenStudy (anonymous):

4. \[\left( \frac{ 6.3\times10^{15} }{ 2.1\times10 ^{11} } \right)\]

OpenStudy (dape):

I'm sure you can do the number 3 you posted, just square everything everything in the parentheses. In number 4 you do like in the first problems, subtract the exponent in the denominator from the one in the numerator (and you also need to calculate 6.3/2.1)

OpenStudy (anonymous):

so number 3 would be \[\frac{ 4 }{ 25n ^{18} }\]

OpenStudy (dape):

Very true :)

OpenStudy (anonymous):

Thanks and number 4 is \[3\times10^{4}\]

OpenStudy (dape):

Yeah, you got it

OpenStudy (anonymous):

oh thank you so much!

OpenStudy (dape):

You're welcome :)